OpenStudy (anonymous):

Solve trig equation 0

OpenStudy (agent0smith):

\[\LARGE \sqrt2 \sin \left( \frac{ x }{ 2} \right) \tan \left( \frac{ x }{ 2} \right) +\sqrt2 \cos \left( \frac{ x }{ 2 } \right) = 2\] ...?

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

For simplicity, i'm replacing all the x/2 with u. So u = x/2, and i'm changing the tan into sin/cos \[\LARGE \sqrt2 \sin u \frac{ \sin u }{\cos u }+\sqrt2 \cos u = 2\] Start by multiplying both sides by cos u. Then you can use the identity sin^2 = 1 - cos^2 and you'll eventually get what looks like a quadratic equation.... i gotta leave to eat, hopefully someone will finish.